OpenStudy (anonymous):

okay i got three points from my equation (-y=-x+1) but they where wrong. I got -1,0 -2, 1 -3,2 can somebody explain why its wrong they make a line on a graph though.

OpenStudy (binary3i):

what are the points? (-1,0) (-2,1) (-3,2) ?

OpenStudy (anonymous):

yeah thats whats i got from it so far.

OpenStudy (binary3i):

put the valus of x in the equation and get your y. eg. put x= -1 -y = -(-1) +1 -y=2 y=-2 point is (-1,-2) you do others.

OpenStudy (anonymous):

okay thanks.